# Challenging Excited States from Adaptive Quantum Eigensolvers: Subspace Expansions vs. State-Averaged Strategies

### v1

@@ 2.2. MORE-ADAPT-VQE
    An objective function is used based on the weighted energies of an arbitrary number of states.
@@
    $\hat{H}$ is diagonalized in the basis of these approximate states to obtain more accurate energies and wavefunctions.
@@
    $g_{j}=\frac{\partial}{\partial\theta_{j}}\sum_{i}\omega_{i}\epsilon_{ij}\bigg{|}_{\theta_{j}=0}$ (4)
@@
    Perform a multistate contracted [32] (i.e. state-averaged) VQE subroutine where $\boldsymbol{\theta}$ is optimized to minimize $E_{\mathrm{SA}}$:
@@
    Compute the elements of the dressed Hamiltonian $\bar{H}$, where
    $\bar{H}_{ij}$ $=\braket{\phi_{i}}{\hat{U}^{\dagger}\left(\boldsymbol{\theta}\right)\hat{H}\hat{U}\left(\boldsymbol{\theta}\right)}{\phi_{j}}$ (8)
    Then diagonalize $\bar{H}$ in the basis of the reference states:
    $\bar{H}\ket{\Psi_{i}}$ $=E_{i}^{\prime}\ket{\Psi_{i}}$ (9)
  7. 7.
    If a convergence criterion has not yet been achieved, return to step 3. Otherwise, return the set of energies $\left\{E_{i}^{\prime}\right\}$. Each $E_{i}^{\prime}$ is bounded below by the true $i$th lowest energy, a useful feature of the Ritz diagonalization [50].
@@
Other operators $\hat{O}$ can also be measured in the basis of $\{\hat{U}\ket{\phi_{i}}\}$, then rotated based on the Ritz eigenvectors of $\hat{H}$ to yield elements $\{\braket{\Psi_{i}}{\hat{O}}{\Psi_{j}}\}$. This is useful in computing other properties of the wavefunction, such as the expectation value of $\hat{S}^{2}$ or the electric dipole operator. We point out that MORE-ADAPT-VQE does not necessarily monotonically improve every energy with every iteration. While the $i^{\mathrm{th}}$ estimate of the energy is bounded below by the true $i^{\mathrm{th}}$ FCI energy [50], MORE-ADAPT-VQE will increase one energy if it means lowering the average energy, as one would expect of a state-averaged method.
@@
We use q-sc-EOM in conjunction with traditional ADAPT-VQE to obtain energies for comparison with MORE-ADAPT-VQE. The expansion manifold of q-sc-EOM is chosen to match that of MORE-ADAPT-VQE in each simulation to achieve a fair comparison. This manifold is not generally optimal, but we aim to demonstrate that MORE-ADAPT-VQE is able to compensate for the deficiences of its references in a way that an expansion-style correction to ADAPT-VQE cannot.
@@ 3. Numerical simulations
where the indices $p,q,r,s$ run over all fermionic modes. In the absence of noise, operators that violate $\hat{S}_{\mathrm{z}}$, particle number, or point group symmetry will never be chosen by ADAPT-VQE for these problems, and can be ignored. The operators do not generally preserve $\hat{S}^{2}$ symmetry, however. All VQE subroutines were performed using the Broyden–Fletcher–Goldfarb–Shanno (BFGS) algorithm to minimize the objective function [56]. FCI and CASSCF results were obtained using Psi4 [57]. Single-reference ADAPT-VQE, MORE-ADAPT-VQE, and q-sc-EOM results were obtained with QForte [58].
@@ 3.1. Rectangular H4 dissociation
We first consider the dissociation of rectangular H4 into two H2 molecules. We targeted six low-lying states in analogy with reference 59. The dissociation trajectory from that study was also used, where the H2 molecules have a fixed bond length of 1 $\mathrm{\AA{}}$. D2h point group symmetry was used. The molecular geometry as a function of the H2-H2 distance $r$ is depicted in Panel A of Figure 1. The canonical Hartree–Fock orbitals at the square geometry are shown in Panel B, corresponding to a 36-determinant Hilbert space. Canonical orbitals were used for all methods in studying this system.
@@
For both quantum excited state approaches, six configuration state functions (CSFs) were used (see Table 1). These were chosen due to their dominant contributions to the eigenstates of interest at the square geometry. For each geometry, we first allowed traditional, single-reference ADAPT-VQE to run on one reference until the FCI energy was recovered, obtaining an ansatz $\hat{U}$ in the process. We then approximated the excited state energies with a q-sc-EOM diagonalization in the subspace of the remaining 5 CSFs. We note that in this case, due to the exactness of the ground state, there is no coupling through $\bar{H}$ between the reference and the other CSFs, so that this is equivalent to simpy diagonalizing in the basis of all six CSFs. Given the degeneracy at 1 Å$1\text{ }\mathrm{\AA{}}$, we report q-sc-EOM results for ADAPT-VQE calculations using both the |$\ket{1\mathrm{a}_{\mathrm{g}}^{2}1\mathrm{b}_{2\mathrm{u}}^{2}}$ and |$\ket{1\mathrm{a}_{\mathrm{g}}^{2}1\mathrm{b}_{3\mathrm{u}}^{2}}$ determinants as references, corresponding to the crossing of the HOMO and LUMO at $r=1$$\mathrm{\AA{}}$. In Panel C of Figure 1, curves from ADAPT-VQE combined with q-sc-EOM are reported, where the determinant of lowest energy is always used as the reference.
@@
We report MORE-ADAPT-VQE results for the rectangular H4 dissociation based on the same set of references shown in Table 1. We did not privilege any of the references. The resulting energy curves are pictured in Panel D of Figure 1. 50 operators were used, by which point all of the FCI energies had been recovered. In obtaining these exact energies, we demonstrate that a fundamental limitation of a single-reference ADAPT-VQE calculation can be addressed by tailoring the ansatz itself to target multiple eigenstates of $\hat{H}$.
This result is fairly surprising. While there necessarily exists a $\hat{U}$ that block-diagonalizes $\hat{H}$ to isolate a block containing the targeted excited state spectrum, it is not obvious that this $\hat{U}$ could be prepared with the UCCGSD ansatz, let alone discovered by a greedy search algorithm like MORE-ADAPT-VQE. We discuss this further in section 4. While H4 is a small system with solutions of mostly different symmetries, we will see excellent agreement between MORE-ADAPT-VQE and FCI for more states with shared symmetries in the case of linear H4, as well as for larger, 7-orbital BeH2 calculations.
@@ 3.2. Linear H4 dissociation
Linear hydrogen chains are commonly used as prototypes for strongly correlated molecules, including for ADAPT-VQE [60, 22]. We consider the specific case of an equally spaced chain of four hydrogens, as shown in Panel A of Figure 2. The system’s canonical Hartree–Fock orbitals at the equilibrium geometry ($r=0.88$$\mathrm{\AA{}}$) are shown in Panel B of Figure 2. The canonical orbitals were used in all methods. For this system, we targeted only Ag states using C2v point group symmetry, giving a Hilbert space of 19 determinants.
@@
For both quantum methods, nine individual Slater determinants of Ag symmetry were used as references based on their large contributions to the exact eigenstates at the equilibrium geometry (0.88 $\mathrm{\AA{}}$): |$\ket{1\mathrm{a}_{\mathrm{g}}^{2}1\mathrm{b}_{2\mathrm{u}}^{2}}$, |$\ket{1\mathrm{a}_{\mathrm{g}}^{2}2\mathrm{a}_{\mathrm{g}}^{2}}$, |$\ket{1\mathrm{b}_{2\mathrm{u}}^{2}2\mathrm{a}_{\mathrm{g}}^{2}}$, |$\ket{1\mathrm{a}_{\mathrm{g}}^{2}1\mathrm{b}_{2\mathrm{u}}^{\alpha}2\mathrm{b}_{2\mathrm{u}}^{\beta}}$, |$\ket{1\mathrm{a}_{\mathrm{g}}^{\alpha}1\mathrm{b}_{2\mathrm{u}}^{2}2\mathrm{a}_{g}^{\beta}}$, and |$\ket{1\mathrm{a}_{\mathrm{g}}^{\alpha}1\mathrm{b}_{2\mathrm{u}}^{\alpha}2\mathrm{a}_{\mathrm{g}}^{\beta}2\mathrm{b}_{2\mathrm{u}}^{\beta}}$, as well as the spin complements of the latter three determinants.
For each geometry, a single-reference ADAPT-VQE calculation was run to completion based on the |$\ket{1\mathrm{a}_{\mathrm{g}}^{2}1\mathrm{b}_{2\mathrm{u}}^{2}}$ determinant. 20 operators were needed to exactly reproduce the FCI ground state energy curve. The remaining 8 determinants were then used as an expansion manifold for q-sc-EOM. The resulting ground and excited states are shown in Panel C of Figure 2.
@@
MORE-ADAPT-VQE calculations were run using all nine determinants as equally weighted references. The resulting energy curves from 100 iterations of MORE-ADAPT-VQE are shown in Panel D of Figure 2, by which point most of the MORE-ADAPT-VQE calculations had obtained the exact FCI energies for all nine states.
@@
Of particular interest is the quintet state. While the reference determinants do not span the full space of $M_{s}=0$ determinants that form the quintet CSFs, MORE-ADAPT-VQE was able to capture this state. Even though the operator pool is not spin symmetry-preserving, traditional ADAPT-VQE has no incentive to break the spin symmetry of the excited states spanned by the references. For states with significant open-shell character with $M_{s}=0$, some important references will consist of linear combinations of many determinants. The higher the number of determinants in the references, the more expensive it becomes to compute off-diagonal elements of $\bar{H}$, resulting in potentially large savings in quantum resources if MORE-ADAPT-VQE is used with single-determinant references.
@@ 3.3. Insertion of beryllium into H2
As a larger model, we consider the insertion of beryllium into an H2 molecule, using a trajectory from reference 62 in the C2v point group. We show the insertion pathway in Panel A of Figure 3. This insertion is a common case study for excited state methods [63, 64]. The S0-S1 avoided crossing at the transition state ($y=1.23$ a0) is particularly challenging to treat as the determinant composition of the S0 state changes along the geometry coordinate. For this reason, in all calculations, molecular orbitals were obtained from a state-specific CASSCF calculation with two electrons distributed in the 1b2 and 3a1 active orbitals. The optimized CASSCF orbitals for the S0 transition state geometry ($y=1.23$ a0) are shown in Panel B of Figure 3. Computations based on quantum methods were performed using these orbitals in the full (6,7) Hilbert space of 1225 determinants.
@@
To obtain the four energy curves for the BeH2 insertion, four reference CSFs were used, shown in Table 2. These CSFs were chosen based on their large contributions to the FCI wavefunctions at the transition state geometry. A minor technical complication must be addressed for the linear ($y=2.54$) a0 geometry. In the linear molecule, the 2px and 2pz orbitals of beryllium can no longer mix with any others. Consequently, CSF 2 is no longer of the same irreducible representation as the desired excited state singlet, and was replaced by a B1u CSF denoted as $2^{*}$, where D2h symmetry was utilized. This also applies to the MORE-ADAPT-VQE simulations.
@@
A singlet ground state energy curve was obtained by running traditional ADAPT-VQE until convergence based on CSF 1 from Table 2. ADAPT-VQE was able to obtain the exact ground state for all points with 225 operators. The other three CSFs were then used as an expansion basis for q-sc-EOM. The resulting curves are shown in Panel C of Figure 3.
@@
Of more interest is the singlet-singlet transition. In Figure 4, we see three key problems with the calculation of this transition dipole moment with q-sc-EOM. The most obvious issue, consistent with the severe problems in computing the energy of the S1 state in Figure 3, is the incorrect behavior of the method as the molecule becomes more linear. Here, q-sc-EOM fails to recognize a region of very large transition dipole moments due to a poor description of the S1 state. As mentioned previously, the references were chosen to represent the excited states of BeH2 at the avoided crossing. Due to ADAPT-VQE’s agnosticism toward the excited state energies, q-sc-EOM fails to obtain good results when the reference manifold becomes less appropriate, even when allowed to run to completion. A similar problem occurs on the left-hand side of the curve, where there is an avoided crossing between S1 and S2, resulting in a relatively small change in the energies but a large change the wavefunctions. This is an illustrative example. While ADAPT-VQE does not give perfectly smooth curves due to the complicated, discrete optimization involved, it makes intuitive sense that for a ground state that does not rapidly change, the unitary $\hat{U}$ should be a slowly varying function of the geometry. Given that the same references are used, this should result in an effective Hamiltonian that does not undergo a rapid change during an avoided crossing between excited states. Because S2 is not well-represented in the effective Hamiltonian throughout the rest of the curve, one would not expect it to abruptly become well-represented just because it becomes lower in energy than S1. This problem is mitigated by MORE-ADAPT-VQE, which is incentivized to rapidly change the character of its excited state solutions, even when the ground state solution does not change. We finally note that, even ignoring the jagged curve, the peak transition dipole moment about the transition state geometry is significantly overestimated by q-sc-EOM, an unsurprising result due to the inadequate characterization of S1.

### v2

@@ 2.2. MORE-ADAPT-VQE
    An objective function is used based on the weighted energies of an arbitrary number of states. This feature is common to the subspace-search (SS-) [31] and multistate contracted (MC-) [32] VQEs.
@@
    $\hat{H}$ is diagonalized in the basis of these approximate states to obtain more accurate energies and wavefunctions. This diagonalization scheme is only used in MC-VQE.
@@
    $g_{j}=\frac{\partial}{\partial\theta_{j}}\sum_{i}\omega_{i}\epsilon_{ij}\bigg{|}_{\theta_{j}=0}.$ (4)
@@
    Perform an MC- (i.e. state-averaged) VQE subroutine where $\boldsymbol{\theta}$ is optimized to minimize $E_{\mathrm{SA}}$:
@@
    Then, as in MC-VQE, compute the elements of the dressed Hamiltonian $\bar{H}$, where
    $\bar{H}_{ij}$ $=\braket{\phi_{i}}{\hat{U}^{\dagger}\left(\boldsymbol{\theta}\right)\hat{H}\hat{U}\left(\boldsymbol{\theta}\right)}{\phi_{j}}.$ (8)
    Then diagonalize $\bar{H}$ in the basis of the reference states {|$\left\{\ket{\phi_{i}}\right\}$ to obtain the eigenstates {|$\left\{\ket{\psi_{i}}\right\}$:
    $\bar{H}\ket{\psi_{i}}$ $=E_{i}^{\prime}\ket{\psi_{i}}$ (9)
    The resulting eigenstates will be linear combinations of the references with coefficients $\left\{c_{ij}\right\}$:
    |$\ket{\psi_{i}}=\sum_{j}c_{ji}\ket{\phi_{j}}$ (10)
    We will define
    |$\ket{\Psi_{i}}=\hat{U}\ket{\psi_{i}},$ (11)
    such that |$\ket{\Psi_{i}}$ is the actual approximate eigenstate of the original Hamiltonian $\hat{H}$.
@@
    If a convergence criterion has not yet been satisfied, return to step 3. Otherwise, return the set of energies $\left\{E_{i}^{\prime}\right\}$. Each $E_{i}^{\prime}$ is bounded below by the true $i$th lowest energy, a useful feature of the Ritz diagonalization [50].
@@
Other operators $\hat{O}$ can also be measured in the basis of $\{\hat{U}\ket{\phi_{i}}\}$, then rotated based on the Ritz eigenvectors of $\hat{H}$ to yield elements $\{\braket{\Psi_{i}}{\hat{O}}{\Psi_{j}}\}$. This is useful in computing other properties of the wavefunction, such as the expectation value of $\hat{S}^{2}$ or the electric dipole operator. We point out that MORE-ADAPT-VQE does not necessarily monotonically improve every energy with every iteration. While the $i^{\mathrm{th}}$ estimate of the energy is bounded below by the true $i^{\mathrm{th}}$ FCI energy [50], MORE-ADAPT-VQE will increase one energy if it means lowering the average energy, as one would expect of a state-averaged method. A real implementation of MORE-ADAPT-VQE would obtain energy and gradient values in steps 3 and 5 by preparing and measuring states on hardware, along with computing the elements of $\bar{H}$. In this work, however, these measurements are simulated in the absence of noise or finite shot count error.
@@
We use q-sc-EOM in conjunction with traditional ADAPT-VQE to obtain energies for comparison with MORE-ADAPT-VQE. The expansion manifold of q-sc-EOM is chosen to match that of MORE-ADAPT-VQE in each simulation to achieve a fair comparison , as opposed to a large manifold based on a maximum excitation rank, such as the more commonly used set of singly and doubly excited determinants. Our chosen manifolds are not generally optimal, but we aim to demonstrate that MORE-ADAPT-VQE is able to compensate for the deficiences of its references in a way that an expansion-style correction to ADAPT-VQE cannot.
@@ 3. Numerical simulations
where the indices $p,q,r,s$ run over all fermionic modes. In the absence of noise, operators that violate $\hat{S}_{\mathrm{z}}$, particle number, or point group symmetry will never be chosen by ADAPT-VQE for these problems, and are excluded. The operators do not generally preserve $\hat{S}^{2}$ symmetry, however. In all MORE-ADAPT-VQE calculations, equal weights $\frac{1}{k}$ were used for each reference. We discuss the implications of this choice in section 4. All VQE subroutines were performed using the Broyden–Fletcher–Goldfarb–Shanno (BFGS) algorithm to minimize the objective function [56]. FCI and CASSCF results were obtained using Psi4 [57]. All single-reference ADAPT-VQE calculations were allowed to achieve FCI accuracy. In a real application, the exactness of the energy is not a valid stopping criterion, since the FCI energy cannot be known a priori. The use of an exact ground state, however, allows us to isolate the behavior of q-sc-EOM from the problem of an inaccurate ADAPT-VQE calculation. In MORE-ADAPT-VQE calculations, we report the results obtained by terminating the algorithm at specific numbers of operators being included in the ansatz, specified throughout this section. This is similar in spirit to using a circuit depth stopping criterion.
Single-reference ADAPT-VQE, MORE-ADAPT-VQE, and q-sc-EOM results were obtained with QForte [58].
@@ 3.1. Rectangular H4 dissociation
We first consider the dissociation of rectangular H4 into two H2 molecules. We targeted six low-lying states in analogy with reference 59. The dissociation trajectory from that study was also used, where the H2 molecules have a fixed bond length of 1 $\mathrm{\AA{}}$. D2h point group symmetry was used. The molecular geometry as a function of the H2-H2 distance $r$ is depicted in Panel A of Figure 1. The canonical Hartree–Fock orbitals at the square geometry are shown in Panel B. This corresponds to Ag, B1g, B2u, and B3u spaces of 12, 8, 8, and 8 determinanants, respectively, for a total relevant Hilbert space of dimension 36. Canonical orbitals were used for all methods in studying this system.
@@
For both quantum excited state approaches, six configuration state functions (CSFs) were used as the references {|$\left\{\ket{\phi_{i}}\right\}$ (see Table 1). These were chosen due to their dominant contributions to the eigenstates of interest at the square geometry. For each geometry, we first allowed traditional, single-reference ADAPT-VQE to run on one reference until the FCI energy was recovered (error less than $1\times 10^{-13}$ $E_{\mathrm{h}}$), obtaining an ansatz $\hat{U}$ in the process. This required 11 ADAPT-VQE iterations. We then approximated the excited state energies with a q-sc-EOM diagonalization in the subspace of the remaining 5 CSFs. We note that in this case, due to the exactness of the ground state, there is no coupling through $\bar{H}$ between the reference and the other CSFs, so that this is equivalent to simpy diagonalizing in the basis of all six CSFs. Given the degeneracy at 1 Å$1\text{ }\mathrm{\AA{}}$, we report q-sc-EOM results for ADAPT-VQE calculations using both the |$\ket{1\mathrm{a}_{\mathrm{g}}^{2}1\mathrm{b}_{2\mathrm{u}}^{2}}$ and |$\ket{1\mathrm{a}_{\mathrm{g}}^{2}1\mathrm{b}_{3\mathrm{u}}^{2}}$ determinants as references, corresponding to the crossing of the HOMO and LUMO at $r=1$$\mathrm{\AA{}}$. In Panel C of Figure 1, curves from ADAPT-VQE combined with q-sc-EOM are reported, where the determinant of lowest energy is always used as the reference.
@@
We report MORE-ADAPT-VQE results for the rectangular H4 dissociation based on the same set of references shown in Table 1. We did not privilege any of the references. The resulting energy curves are pictured in Panel D of Figure 1. 50 operators were used, by which point all of the FCI energies had been recovered (error less than $1\times 10^{-10}$ $E_{\mathrm{h}}$ for each state). In obtaining these exact energies, we demonstrate that a fundamental limitation of a single-reference ADAPT-VQE calculation can be addressed by tailoring the ansatz itself to target multiple eigenstates of $\hat{H}$.
This result is fairly surprising. While there necessarily exists a $\hat{U}$ that block-diagonalizes $\hat{H}$ to isolate a block containing the targeted excited state spectrum, it is not obvious that this $\hat{U}$ could be prepared with the UCCGSD ansatz, let alone discovered by a greedy search algorithm like MORE-ADAPT-VQE. We discuss this further in section 4. While H4 is a small system with solutions of mostly different symmetries, we also see excellent agreement between MORE-ADAPT-VQE and FCI for more states with shared symmetries in the case of linear H4, as well as for larger, 7-orbital BeH2 calculations.
@@ 3.2. Linear H4 dissociation
Linear hydrogen chains are commonly used as prototypes for strongly correlated molecules, including for ADAPT-VQE [60, 22]. We consider the specific case of an equally spaced chain of four hydrogens, as shown in Panel A of Figure 2. The system’s canonical Hartree–Fock orbitals at the equilibrium geometry ($r=0.88$$\mathrm{\AA{}}$) are shown in Panel B of Figure 2. The canonical orbitals were used in all methods. For this system, we targeted only Ag states using D2h point group symmetry, giving a Hilbert space of 20 determinants.
@@
For both quantum methods, nine individual Slater determinants of Ag symmetry were used as references {|$\left\{\ket{\phi_{i}}\right\}$ based on their large contributions to the exact eigenstates at the equilibrium geometry (0.88 $\mathrm{\AA{}}$): |$\ket{1\mathrm{a}_{\mathrm{g}}^{2}1\mathrm{b}_{2\mathrm{u}}^{2}}$, |$\ket{1\mathrm{a}_{\mathrm{g}}^{2}2\mathrm{a}_{\mathrm{g}}^{2}}$, |$\ket{1\mathrm{b}_{2\mathrm{u}}^{2}2\mathrm{a}_{\mathrm{g}}^{2}}$, |$\ket{1\mathrm{a}_{\mathrm{g}}^{2}1\mathrm{b}_{2\mathrm{u}}^{\alpha}2\mathrm{b}_{2\mathrm{u}}^{\beta}}$, |$\ket{1\mathrm{a}_{\mathrm{g}}^{\alpha}1\mathrm{b}_{2\mathrm{u}}^{2}2\mathrm{a}_{g}^{\beta}}$, and |$\ket{1\mathrm{a}_{\mathrm{g}}^{\alpha}1\mathrm{b}_{2\mathrm{u}}^{\alpha}2\mathrm{a}_{\mathrm{g}}^{\beta}2\mathrm{b}_{2\mathrm{u}}^{\beta}}$, as well as the spin complements of the latter three determinants.
For each geometry, a single-reference ADAPT-VQE calculation was run to completion based on the |$\ket{1\mathrm{a}_{\mathrm{g}}^{2}1\mathrm{b}_{2\mathrm{u}}^{2}}$ determinant (error less than $1\times 10^{-8}$ $E_{\mathrm{h}}$). 20 operators were needed to exactly reproduce the FCI ground state energy curve. The remaining 8 determinants were then used as an expansion manifold for q-sc-EOM. The resulting ground and excited states are shown in Panel C of Figure 2.
@@
MORE-ADAPT-VQE calculations were run using all nine determinants as equally weighted references. The resulting energy curves from 100 iterations of MORE-ADAPT-VQE are shown in Panel D of Figure 2, by which point most of the MORE-ADAPT-VQE calculations had obtained the exact FCI energies for all nine states (error less than $1\times 10^{-10}$ $E_{\mathrm{h}}$ for each state).
@@
Of particular interest is the quintet state. While the reference determinants do not span the full space of $M_{s}=0$ determinants that form the quintet CSFs, MORE-ADAPT-VQE was able to capture this state. Even though the operator pool is not spin symmetry-preserving, traditional ADAPT-VQE has no incentive to break the spin symmetry of the excited states spanned by the references. For states with significant open-shell character with $M_{s}=0$, important CSFs will consist of linear combinations of many determinants. The greater the number of determinants in the references, the more expensive it becomes to compute off-diagonal elements of $\bar{H}$, resulting in potentially large savings in quantum resources if MORE-ADAPT-VQE is used with single-determinant references. Given the severity of spin-contamination in the q-sc-EOM results for linear H4 relative to the q-sc-EOM results for the other two systems, which used CSFs instead, using individual determinants is a poor idea for small-manifold q-sc-EOM. The comparatively forgiving nature of MORE-ADAPT-VQE in dealing with spin-contaminated references is a very attractive feature.
@@ 3.3. Insertion of beryllium into H2
As a larger model, we consider the insertion of beryllium into an H2 molecule, using a trajectory from reference 62 in the C2v point group. We show the insertion pathway in Panel A of Figure 3. This insertion is a common case study for excited state methods [63, 64]. The S0-S1 avoided crossing at the transition state ($y=1.23$ a0) is particularly challenging to treat as the determinant composition of the S0 state changes along the geometry coordinate. For this reason, in all calculations, molecular orbitals were obtained from a state-specific CASSCF calculation with two electrons distributed in the 1b2 and 3a1 active orbitals. The optimized CASSCF orbitals for the S0 transition state geometry ($y=1.23$ a0) are shown in Panel B of Figure 3. Computations based on quantum methods were performed using these orbitals in the full (6,7) active space.
The individual A1 and B2 subspaces contain 321 and 304 determinants, respectively, for a total relevant Hilbert space of 625 determinants. Note that CASSCF is used only to choose a better set of orbitals prior to the quantum algorithms. The orbitals are fixed throughout MORE-ADAPT-VQE, unlike in reference [34].
@@
To obtain the four energy curves for the BeH2 insertion, four CSFs were used as the references {|$\left\{\ket{\phi_{i}}\right\}$, shown in Table 2. These CSFs were chosen based on their large contributions to the FCI wavefunctions at the transition state geometry. A minor technical complication must be addressed for the linear ($y=2.54$ a0) geometry. In the linear molecule, the 2px and 2pz orbitals of beryllium can no longer mix with any others. Consequently, CSF 2 is no longer of the same irreducible representation as the desired excited state singlet, and was replaced by a B1u CSF denoted as $2^{*}$, where D2h symmetry was utilized. This also applies to the MORE-ADAPT-VQE simulations. This results in solutions of Ag, B1u, and B3g symmetry, with Hilbert space dimensions of 169, 152, and 148, respectively, for a total relevant Hilbert space dimension of 469.
@@
A singlet ground state energy curve was obtained by running traditional ADAPT-VQE until convergence (error less than $1\times 10^{-7}$ $E_{\mathrm{h}}$) based on CSF 1 from Table 2. ADAPT-VQE was able to obtain the exact ground state for all points with 225 operators. The other three CSFs were then used as an expansion basis for q-sc-EOM. The resulting curves are shown in Panel C of Figure 3.
@@
Of more interest is the singlet-singlet transition. In Figure 4, we see three key problems with the calculation of this transition dipole moment with q-sc-EOM. The most obvious issue, consistent with the severe problems in computing the energy of the S1 state in Figure 3, is the incorrect behavior of the method as the molecule becomes more linear. Here, q-sc-EOM fails to recognize a region of very large transition dipole moments due to a poor description of the S1 state. As mentioned previously, the references were chosen to represent the excited states of BeH2 at the avoided crossing. Due to ADAPT-VQE’s agnosticism toward the excited state energies, q-sc-EOM fails to obtain good results when the reference manifold becomes less appropriate, even when allowed to run to completion. A similar problem occurs on the left-hand side of the curve, where there is an avoided crossing between S1 and S2, resulting in a relatively small change in the energies but a large change in the wavefunctions. This is an illustrative example. While ADAPT-VQE does not give perfectly smooth curves due to the complicated, discrete optimization involved, it makes intuitive sense that for a ground state that does not rapidly change, the unitary $\hat{U}$ should be a slowly varying function of the geometry. Given that the same references are used, this should result in an effective Hamiltonian that does not undergo a rapid change during an avoided crossing between excited states. Because S2 is not well-represented in the effective Hamiltonian throughout the rest of the curve, one would not expect it to abruptly become well-represented just because it becomes lower in energy than S1. This problem is mitigated by MORE-ADAPT-VQE, which is incentivized to rapidly change the character of its excited state solutions, even when the ground state solution does not change. We finally note that, even ignoring the jagged curve, the peak transition dipole moment about the transition state geometry is significantly overestimated by q-sc-EOM, an unsurprising result due to the inadequate characterization of S1.
@@
For completeness, we include in Figure 5 some information about the individual ansätze constructed for each system at each geometry. We note a few interesting trends. First, the number of traditional ADAPT-VQE iterations seems to be approximately bounded by the dimension of the size of the symmetric irreducible representation, though for the larger BeH2 system, far fewer operators are required. MORE-ADAPT-VQE needs an ansatz $\hat{U}$ which rotates $k$ references into $k$ FCI wavefunctions in Hilbert spaces of sizes $|\mathcal{H}_{1}|,|\mathcal{H}_{2}|,\dots,\text{ and }|\mathcal{H}_{k}|$, respectively. One might expect that the parameters required for MORE-ADAPT-VQE would be approximately bounded by the sum of the Hilbert space dimensions. For example, the rectangular H4 calculations in this work would correspond to
$\sum_{i}|\mathcal{H}_{i}|=12+12+8+8+8+8=56\text{ determinants}.$ (16)
MORE-ADAPT-VQE, using only 50 operators, is within this bound. Similarly, linear H4 would be expected to require roughly 180 operators for its 9 totally symmetric solutions. This is far larger than the 100 operators which give all 9 exact FCI states for most bond distances. This suggests that for large numbers of references with the same irreducible representation, the number of required operators scales significantly better than simply multiplying the size of the irreducible representation by $k$. If one is not concerned with balanced treatment of solutions of different irreducible representations, the set of references can always be divided up into separate MORE-ADAPT-VQE calculations of distinct irreducible representations. Further investigation of this idea will be reserved for future work. Based on the proposed linear upper bound, BeH2 would be expected to require 1267 operators (aside from the linear geometry at $y=2.54$ a0). We see that excellent results are obtained with only 300 operators, indicating that the large upper bound will be a gross over-estimation in larger systems. With respect to the additional operators used by MORE-ADAPT-VQE, it is clear that a larger fraction of the operator pool must be used relative to traditional ADAPT-VQE. That is, the MORE-ADAPT-VQE ansätze are not simply applying the same operators additional times, even in the case of linear H4 where the excited states are of the same spatial symmetry as the ground state. These results are encouraging and suggest that further studies of larger systems are needed to determine the general viability of the MORE-ADAPT-VQE method in terms of ansatz length.
Figure 5: Ansätze information from numerical simulations of (A) rectangular H4 dissociation, (B) linear H4 dissociation, and (C) BeH2 insertion. The geometry coordinates are the same as those in Figures 1-3. The dashed black lines correspond to the totally symmetric FCI Hilbert space dimension associated with the traditional ADAPT-VQE solution, while the solid black lines correspond to the FCI Hilbert space containing all irreducible representations associated with the set of target states for each system. Note that in the case of linear H4, only Ag states were targeted, so that the two FCI spaces are identical. The green lines correspond to the number of operators in the UCCGSD pool for each system. Recall that the UCCGSD pool does not include operators which are not totally symmetric, since they cannot be chosen by the ADAPT-VQE procedure. It follows that rectangular and linear H4 have different UCCGSD pools. Note that both the FCI spaces and the operator pool decrease in size as BeH2 becomes linear at $y=2.54$$\mathrm{\AA{}}$ in panel (C), as indicated by the grey shaded region. At that point, the target space includes the Ag, B1u, and B3u irreducible representations. The red and blue lines correspond to the single-reference ADAPT-VQE calculation needed for q-sc-EOM and the MORE-ADAPT-VQE calculation, respectively. The solid red and blue lines indicate the total number of operators used in each ansatz, while the dashed lines indicate the number of distinct operators used.